OpenStudy (anonymous):

person kicks ball directed 53 degree above horizontal .ball lands on roof 7.2 m high.wall of building is 25 m away,and it takes ball 2.1s to pass directly over wall. calculate a)initial velocity of ball b)horizontal range of ball c)by what vertical distance does ball clear wall of building?

OpenStudy (anonymous):



OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

yup..

OpenStudy (anonymous):



OpenStudy (anonymous):

what's that vector (black, horizontal) in terms of the angle and V?

OpenStudy (anonymous):

V is the red vector

OpenStudy (anonymous):

\[v_{x}..\]

OpenStudy (anonymous):

Vcos(53) is vx 25m = Vcos(53) * t

OpenStudy (anonymous):

"it takes ball 2.1s to pass directly over wall" 25m = Vcos(53) * (2.1)

OpenStudy (anonymous):

so V=?

OpenStudy (anonymous):

vx=11.9m/s

OpenStudy (anonymous):

check again...

OpenStudy (anonymous):

divide both sides by 2.1 ... divide both sides by cos(53)...

OpenStudy (anonymous):

vx=vcos theta 11.9=v cos53 v=20 m/s

OpenStudy (anonymous):

this is initial velocity of balll right??

OpenStudy (anonymous):

19.8 ok close enough I guess...

OpenStudy (anonymous):

so that's a) b)...

OpenStudy (anonymous):

now total horizontal range dx...???

OpenStudy (anonymous):

ball starts at y=0 ends at y=7.2m (on the roof) so: 7.2 = Vsin(53) *t - 1/2 * g * t^2

OpenStudy (anonymous):

if you can find that t, you can find horizontal range... (because we already know horizontal velocity)

OpenStudy (anonymous):

are you able to solve that for t? (hint: it's a quadratic equation)

OpenStudy (anonymous):

m tryin..1 sec..

OpenStudy (anonymous):

set it equal to zero a = -4.9 b= 15.8 c= -7.2

OpenStudy (anonymous):

approx.3.67 s=4s

OpenStudy (anonymous):

2.67s

OpenStudy (anonymous):

i took a=4.9,b=-16,c=-7.2

OpenStudy (anonymous):

b^2 = 250 -4ac = -141.1 b^2 -4ac = 108.9 sqrt(b^2 -4ac) = 10.4 -b -sqrt(b^2 -4ac) = -15.8 -10.4 =-26.2 2a = - 9.8 -26.2/-9.8 = 2.67s

OpenStudy (anonymous):

sign errors reposted.

OpenStudy (anonymous):

so now you can find horizontal range range = 11.9*2.67

OpenStudy (anonymous):

your c should be+7.2 ryt??

OpenStudy (anonymous):

-7.2

OpenStudy (anonymous):

7.2 =Vsin(53) *t - 1/2 * g * t^2

OpenStudy (anonymous):

set equal to zero 0 = -7.2 +Vsin(53) *t - 1/2 * g * t^2

OpenStudy (anonymous):

ok now i got my mistake..sorry

OpenStudy (anonymous):

cool.